OpenStudy (anonymous):

hw do u solve 6-p under square root sign =p?

OpenStudy (anonymous):

you mean√6-p = p. so 6-p=p^2.(p≤6) p^2+p-6=0.(p≤6) (p+3)(p-2)=0.(p≤6) p1=-3,p2=2.

OpenStudy (anonymous):

no the -p is also under the square root sign

OpenStudy (anonymous):

so √6-p = √p , is that what you mean?

OpenStudy (anonymous):

i've got it right . √(6-p) = p. so 6-p=p^2.(p≤6) p^2+p-6=0.(p≤6) (p+3)(p-2)=0.(p≤6) p1=-3,p2=2.

OpenStudy (anonymous):

no like the square root sign over 6-p and =p out of it

OpenStudy (anonymous):

yes, the answer is correct, p= -3 or 2.